Help please !
Find the area of the shaded region. The graph to the right depicts IQ scones of aduits, and hose scores are nomaly of 15 The area of the shaded region is□(Round to four decimal places as needed)
Answer #1

Solution:- The are of the shaded region is 0.3662.

Mean = 100, S.D = 15

x1 = 102

x2 = 121

By applying normal distribution:-

z = \frac{x-\mu }{\sigma }

z1 = 0.1333

z2 = 1.40

P( 0.1333 < z < 1.40) = P(z > 0.1333) - P(z > 1.40)

P( 0.1333 < z < 1.40) = 0.447 - 0.0808

P( 0.1333 < z < 1.40) = 0.3662